OpenStudy (anonymous):

examine the following sets are linear independence or dependence: u1=e^ix, u2=e^-ix, u3= sinx.

OpenStudy (anonymous):

u1= e^ix = Cos(x) + i Sin(x) u2= e^-ix= Cos(-x) + i Sin(-x) = Cos(x) - i Sin(x) u3= Sin(x) they are linearly dependent since you are able to write u1 - u2 =2i u3 theorem: a set of vectors is linearly independent if and only if none of the vectors can be written as a linear combination of the others.

OpenStudy (anonymous):

thank you